# Inside the black box: Neural network-based real-time prediction of US recessions

### v2

@@ 3.3 Performance evaluation
The estimated models produce two types of forecasts, probability and point predictions. The models first generate probability predictions. In the next step, these probabilities are transformed into binary point predictions using a specified threshold. While the second type may not always be necessary, it is included to ensure the forecast is more easily interpretable for the end-users. The performance of both the probability and point predictions is evaluated using various statistical measures. With the exception of metrics specifically related to probability predictions, the evaluation is based on the output of a contingency table, commonly known as a confusion matrix as illustrated in Table 4. This matrix provides a framework for assessing the accuracy of the point predictions.
@@
The table contains elements representing the count of observations belonging to each category. True positives (TP) and true negatives (TN) indicate the correct classification of positive and negative outcomes, respectively. False positives (FP) occur when the prediction is positive, but the actual value is negative. Conversely, false negatives (FN) arise when the prediction is negative, but the actual value is positive.
The confusion matrix is composed of elements that represent the count of observations belonging to different categories. True positives (TP) and true negatives (TN) indicate the correct classification of positive and negative outcomes, respectively. False positives (FP) occur when the prediction is positive while the actual value is negative. Conversely, false negatives (FN) arise when the prediction is negative while the actual value is positive. By utilizing these counts, various performance evaluation measures are derived to assess the overall performance of the models.
@@
Table 5 reports the list of the metrics for the performance evaluation. A carefully selected set of statistical metrics is used to assess the performance from different perspectives. The metrics, Area Under the Receiver Operating Characteristic Curve (AUROC) and Area Under the Precision-Recall Curve (AUPRC) are employed to measure the models’ raw performance in the sense that these metrics do not require any threshold to convert probabilities into binary outcomes, making them suitable for evaluating the models’ pure performance. The Receiver Operating Characteristics (ROC) curve displays the entire set of possible combinations of true positive rates $TPR(c)=\frac{TP(c)}{\#\>Actual\>Positive}$ and false positive rates $FPR(c)=\frac{FP(c)}{\#\>Actual\>Negative}$ for some cutpoint $c\in(0,1)$ that maps the predicted probability to a binary category. Similarly, the Precision Recall (PR) curve plots the complete set of possible combinations of precision and recall for $c\in(0,1)$, where recall is a synonym for sensitivity. Both the area under the ROC curve and the PR curve increase with the underlying metrics for a given cutpoint. By aggregating over the entire set of cutpoints, these curves deliver a framework to assess the pure predictive ability of a forecasting model. Tharwat (2021) provides a comprehensive overview of other metrics, while chapter 2 briefly discusses their strengths and weaknesses, particularly in the context of imbalanced binary classification problems.
@@ 4 Empirical results
Table 6 presents the out-of-sample forecast performance of the models for various forecast horizons. The performance values of the logit and ridge logit models, obtained from chapter 2, are included in the analysis after applying a fixed threshold of 0.5, and can now be directly compared to the neural network models in this study since they are based on the same data and forecast horizons.
@@
The table reports the performance evaluation measures of forecasts obtained by logit models, ridge logit models, feed-forward (FFN), long short-term memory (LSTM), and gated recurrent unit (GRU) neural networks over different forecast horizons for the out-of-sample period, November 2006 to October 2021: Panel (A) presents the nowcasts. Panel (B), (C), (D), and (E) display the 1-months-ahead forecasts, the 3-months-ahead forecasts, the 6-months-ahead forecasts, and the 12-month-ahead forecasts, respectively.
@@
In the nowcasting setup, there is a difference in predictive performance between the types of logit models and neural network models. The standard logit model is not outperforming other models in any considered metric, although the difference is smaller compared to other forecast horizons. The ridge logit model performs similarly to FFN. Although FFN performs slightly better, the differences are small for most of the metrics. However, LSTM and GRU show significantly better performance in terms of MCC and $F_{1}$-Score, mainly due to their high precision values. For instance, LSTM correctly identifies 90% of recession months, while 75% of its positive predictions are correct. In contrast, the ridge logit model has a precision of 51.5%, indicating that LSTM accurately predicts recessions without generating too many false alarms. The AUPRC values also highlight the superiority of LSTM and GRU models over the others, with GRU achieving an AUPRC of 0.837 compared to 0.642 for FFN and 0.529 for the ridge logit model. Davis and Goadrich (2006) argue that PR curves give a more informative picture of an algorithm’s performance, when dealing with imbalanced dataset. In fact, the ratios of the number of months in booms to the number of months in recessions in 194 real-time vintages range from 4.5 to 6.7 with a mean value of 6, which means that the period of booms is in average 6 times longer than that of recessions. The percentage of recessions in the datasets fluctuates around 0.2 across the vintages. Although not extremely skewed, the datasets are certainly imbalanced, and the difference in AUPRC values suggests that LSTM and GRU are better at extracting hidden patterns from scarce data of recessions.
@@
Moving to the immediate-term setup, logit models perform significantly worse, with MCC and $F_{1}$-Score approximately 40% lower than before. Conversely, the other models show similar or slightly improved performance. Among them, GRU exhibits the highest performance across various summarizing metrics such as balanced accuracy, MCC, and $F_{1}$-Score. In the short-term setup, although the ridge logit model demonstrates better predictive performance than LSTM, on average, neural network models, particularly GRU, outperform the other models. GRU maintains a 90% accuracy in classifying recession months with higher specificity and precision, resulting in higher values of balanced accuracy, MCC, and $F_{1}$-Score. In the medium-term setup, LSTM performs slightly better than the other models, while GRU performs poorly, and thus the difference between the two groups is minimal. However, LSTM achieves 90% sensitivity and over 90% specificity, while maintaining precision above 0.5. The overall average performance of the models is the lowest among different forecast horizons. In the long-term setup, GRU surpasses the other models by a significant margin, with an $F_{1}$-Score of 0.737 compared to 0.468 for LSTM and 0.357 for the ridge logit model. GRU still accurately predicts 70% of recession months, with 77.8% of its positive predictions being correct.
@@
The figure depicts precision recall (PR) curves for the out-of-sample period from November 2006 to October 2021: Panel (A) shows the PR curves for the nowcasting forecast horizon, comparing the logit model, ridge logit model, FFN, LSTM, and GRU. Panels (B), (C), (D), and (E) present the PR curves for the immediate-term (1-months-ahead), short-term (3-months-ahead), medium-term (6-months-ahead), and long-term (12-months-ahead) forecast horizon, respectively.
Neural network models, especially LSTM and GRU, show a significant improvement in predictive performance compared to logit models. Across all metrics considered, neural network models consistently outperform logit models by a significant margin, except for the nowcasting setup. This superiority is evident in both pure performance metrics such as AUROC and AUPRC, as well as threshold-based point prediction metrics. The disparity is particularly prominent in AUPRC, which is a more reliable indicator when dealing with imbalanced data, highlighting the neural network models’ ability to handle class imbalance effectively. The performance difference decreases when comparing them to ridge logit models. Ridge logit models exhibit higher AUROC values compared to other model specifications, except for the long-term setup. However, in terms of AUPRC, neural networks again prove to be a better choice. On average, neural network models perform better than ridge logit models in terms of these metrics. The gap widens further when considering only the two recurrent neural network models. Depending on the forecast horizons, either LSTM or GRU may lead the competition, with occasionally substantial differences between them and the other models.
The ROC and PR curves of the models for five different forecast horizons are presented in Figure 6 and 7, respectively. Panel B and C of the ROC curves highlight the noticeable difference in forecast performance between the logit model (represented by the black line) and the other models. In Panel D and E, although the lines are closer together, there is still some discernible gap between each of the models. Overall, the ridge logit model, LSTM, and GRU exhibit superior performance of probability predictions across various forecast horizons. In Figure 7 demonstrates that, the green, blue, and purple lines lie, on average, above the red line and, even more so, above the black line. This finding supports the argument that neural network models, compared to linear models, are more capable of handling class imbalance in the data.
@@
The figure presents out-of-sample recession probabilities for the period from November 2006 to October 2021:: Panel (A) displays the predicted recession probabilities for the nowcasting forecast horizon, comparing the logit model, ridge logit model, FFN, LSTM, and GRU. Panels (B), (C), (D), and (E) show the predicted recession probabilities for the immediate-term (1-months-ahead), short-term (3-months-ahead), medium-term (6-months-ahead), and long-term (12-months-ahead) forecast horizon, respectively. The grey shaded areas depict NBER recession months.
Figure 8 displays the out-of-sample forecasts in probabilities from November 2006 to September 2021, accompanied by the grey shaded areas representing the Great Recession and the Covid-19 recession. Across all panels, similar patterns emerge: the graphs start at high levels before the Great Recession, either remain high or increase during that period, decrease afterward to a range between 0 and 0.25 for an extended period, and then exhibit significant fluctuations around the Covid-19 recession. The models successfully indicate economic and financial downturns during the Great Recession, but an external event like the Covid-19 recession remains largely unpredictable. Depending on the forecast horizon, once the model learns from the impact of the Covid-19 pandemic on the economy, it struggles to interpret this abrupt change and generates alternating forecasts.
@@ 5 Variable importance
To gain a better understanding of the underlying factors driving the results, it is important to identify the main variables that have the most significant impact on the prediction. While linear models can refer to estimated coefficients or marginal effects to measure variable importance, more complex models like neural networks require different approaches. Molnar (2020) provides a comprehensive list of methods that enable the interpretation of complex machine learning models. One such method is Local Interpretable Model-Agnostic Explanations (LIME), which suggests the use of other interpretable models to approximate the predictions locally. For an instance of interest, LIME constructs a new dataset by altering samples variable-wise, drawing from a normal distribution with sample mean and sample standard deviation, and obtains corresponding predictions from the black box model. This modified dataset is then used to train an interpretable model, such as Lasso or a decision tree, which is weighted by the proximity of the sampled instances to the instance of interest. The interpretable model serves as an approximation of the black box model’s predictions at a local level, even though it may not accurately represent the global approximation.
@@
Another solution stems from cooperative game theory, particularly the Shapley value introduced by Shapley (1953). The Shapley value method involves assigning payouts to players based on their contribution to the overall payout. This cooperative framework resembles a game where players form coalitions and receive profits based on their cooperation. In the context of model predictions and their interpretability, the game pertains to the prediction task, and the payout represents the difference between the actual prediction for that data point and the average prediction across all data points. The players correspond to the variable values of the data point, working together to achieve the payout, which means predicting a specific value. To calculate the Shapley value for a specific variable value, all possible coalitions of variable values, excluding the variable of interest, are formed for each data point. The variable values outside a coalition are substituted with random values of those variables from data to generate a prediction. The predictions are computed for each coalition, both with and without the variable value of interest. The difference between these predictions represents the marginal contribution of the variable value for that coalition. This computation is repeated for all possible coalitions, and the average of the marginal contributions across all coalitions yields the Shapley value for that variable value. Finally, the Shapley values for a variable can be averaged across data points to assess the relative importance of variables in comparison to each other.
@@
The Shapley value stands out as the sole explanation method supported by a robust theory that satisfies the axioms of efficiency, symmetry, dummy, and additivity to allow for fair distribution of the contributions among the variables (Molnar (2020)). Conversely, methods like LIME rely on the assumption of local linear behavior of the machine learning model, but lack a theoretical justification for why this approach is effective. Lundberg and Lee (2017) propose an alternative estimation method, the Shapley additive explanations (SHAP), which uses a kernel-based estimation approach for Shapley values (KernelSHAP). SHAP introduces an innovative aspect by presenting the Shapley value explanation in the form of an additive variable attribution method, which can be viewed as a linear model. This perspective establishes a connection between LIME and Shapley values. In this context, SHAP defines the explanation model as follows:
@@
KernelSHAP is a method that estimates the contributions of individual variable values to the predictions by first generating a random coalition $c^{\prime}\in\{0,1\}^{K}$ consisting of K members by flipping a coin multiple times until we obtain a sequence of 0’s and 1’s. The sampled coalition of size $K$ is then used as a data point for the regression model. In this model, the target is the prediction for a coalition, whereby in case of $c_{k}=0$, the absent variable value is substituted with random variable values from the data. This process is repeated for each data point. Then Shapley compliant weights are computed according to the SHAP kernel that is proposed by Lundberg and Lee (2017). Finally, a weighted linear regression model is fitted on the modified data, and the estimated coefficients of the model are the Shapley values.
@@
The main distinction from LIME lies in how instances are weighted in the regression model. In LIME, the weighting is determined based on the proximity of instances to the original instance. The closer an instance is, the higher its weight in LIME. On the other hand, SHAP assigns weights to sampled instances based on the weight they would receive in the Shapley value estimation. Small coalitions (with fewer 1’s) and large coalitions (with many 1’s) receive the highest weights in SHAP. The underlying intuition is that the most knowledge about individual variables can be obtained when their effects can be studied in isolation. For a coalition consisting of a single variable, the isolated main effect of that variable on the prediction can be observed. When a coalition includes all variables except one, it reports the total effect of that particular variable, including both the main effect and variable interactions. However, if a coalition comprises half the variables, it provides limited insight into the contribution of an individual variable due to the numerous possible coalitions with half of the variables.
@@
The SHAP values for logit models are computed by the KernelSHAP method using KernelExplainer from SHAP package in python, whereas the SHAP values for deep learning models are approximated by an enhanced version of the DeepLIFT algorithm introduced by Shrikumar et al. (2017) using DeepExplainer from the same package. DeepLIFT is a method used to analyze the output prediction of a neural network for a specific input. It achieves this by backpropagating the contributions of all units in the network to each input variable. DeepLIFT compares the activation of each unit to its reference activation, where for each variable, its sample mean is used as the reference input, and assigns contribution scores based on the difference between them. Unlike other approaches, DeepLIFT has the ability to uncover dependencies that may be overlooked. It can also consider positive and negative contributions separately. Additionally, the scores can be efficiently computed in a single backward pass.
@@
SHAP method generates a matrix of Shapley values based on the number of data points and variables. These Shapley values can be used to create global explanations by averaging the absolute Shapley values per variable across the twelve time steps and data points. This process is repeated for the entire out-of-sample period. To compare the SHAP results of neural network models to linear models, GRU and the ridge logit model are selected for the subsequent analysis, as they perform better than the others in their respective groups of models. The results for GRUs are presented in Figure 9, which depicts the distribution of average absolute SHAP values across variables. The variables are ranked based on their medians, providing insights into their importance for different forecast horizons.
@@
The ranking of variables may vary over time, but it offers an initial understanding of the key drivers influencing the prediction results for different forecast horizon. Variables associated with financial conditions such as S&P 500 index and term spread, and macroeconomic variables related to GDP, inflation, and house market, which are commonly regarded as important recession indicators, consistently rank high (Estrella and Mishkin (1998)). These findings provide valuable insights into the black box and suggest that neural networks effectively capture the patterns inherent in the dynamics of the frequently cited recession indicators.
@@
Figure 10 presents the SHAP results for ridge logit model. There are three major differences, among others, between the results in Figure 9 and 10. Firstly, the order of variables differs for each forecast horizon between the two types of models. While some variables appear in the leading groups for both models, their specific order varies significantly. For instance, the money supply M2 is influential in the ridge model but less prominent in the GRU model. Spearman’s rank correlations between them range from -0.24 to 0.10, indicating weak correlation.
@@
This suggests that GRU and ridge models assign different weights to the variables. This difference could be attributed to the neural network’s ability to capture nonlinear relationships, allowing for varied weighting of the variables based on different patterns. Secondly, the SHAP values of the GRU models exhibit a more even distribution among the variables compared to the ridge model. This is reflected in the smaller variation in medians between the variables. The distinction becomes particularly pronounced in the short-term setting (up to 3 months), where the ridge model heavily depends on the S&P 500 index. Additionally, the ridge model exhibits numerous outliers in the SHAP values, indicating its high sensitivity to changes in macroeconomic and financial conditions in SHAP value estimation. In contrast, this highlights GRU as a more robust modelling framework. Finally, the within variation of the GRU-based SHAP values for some variables are significantly larger than those of the others. Certain variables demonstrate little change in their SHAP values, while others exhibit significant variation. This phenomenon is less prominent in the outcomes of the ridge model. The variances are more uniformly spread, evident in the roughly equal sizes of boxes across variables. This could be viewed as another indication of the GRU’s more resilient modeling framework. The GRU is capable of adapting to shifts in economic conditions, adjusting the weights of specific variables accordingly. In contrast, the ridge model has relatively fixed variable orders, with limited flexibility for variation.
The marginal effect analysis is another option to evaluate the variable importance of linear models. Since deep learning models can not be interpreted in terms of marginal effects, the average marginal effects of the ridge model are computed for the out-of-sample period and plotted in Figure 11. The Spearman’s rank correlations between the variable series for the ridge model, evaluated using either SHAP values or marginal effects, range from 0.22 to 0.51. The magnitudes of the correlations are not especially large, but the previous findings, especially the first one about the leading recession indicators, also apply to the ridge model evaluated by the marginal effects. The averaged LIME values in Figure 12 which may be regarded as the marginal effect version of GRU reaffirm the main findings discussed above. The only minor distinction is that it places greater emphasis on the term spread as the comprehensive recession indicator, both in short-term and long-term contexts.
@@
The SHAP method allows for a detailed analysis of the impact of a single variable on predictions. In order to anticipate recessions as early as possible, I focus on the long-term forecast horizon (12 months) and consider the term spread as the most relevant and influential variable for GRU to examine its effects on recession predictions for the GRU model. Figure 13 presents the dependence plot of the term spread using the latest available data from October 2021. The x-axis represents the standardized value of the variable (term spread) while the y-axis represents the corresponding SHAP value. Notably, there is a cutoff point around 0.2, above which a higher value of the term spread has a negative impact on the predicted recession probability, while a value below the cutoff positively influences the predicted probability. Additionally, the farther the value deviates from the cutoff, the stronger the effect of the term spread on the predicted probability becomes. This corroborates the traditional interpretation of the term spread in relation to the probability of a recession according to Estrella and Mishkin (1996) and indicates that advanced neural network models like GRU uncovers and utilizes this relationship.

### v3

@@ 3.3 Performance evaluation
The estimated models produce two types of forecasts: probability and point predictions. The models first generate probability predictions. The next step transforms these probabilities into binary point predictions using a specified threshold. While the second type may not always be necessary, it is included to ensure the forecast is more easily interpretable for the end users. The performance of both the probability and point predictions is evaluated using various statistical measures. Except for metrics related explicitly to probability predictions, the evaluation is based on the output of a contingency table, commonly known as a confusion matrix, as illustrated in Table 4. This matrix provides a framework for assessing the accuracy of the point predictions.
@@
The table contains elements representing the count of observations belonging to each category. True positives (TP) and true negatives (TN) indicate the correct classification of positive and negative outcomes. False positives (FP) occur when the prediction is positive, but the actual value is negative. Conversely, false negatives (FN) arise when the prediction is negative, but the actual value is positive.
The confusion matrix comprises elements representing the count of observations belonging to different categories. True positives (TP) and true negatives (TN) indicate the correct classification of positive and negative outcomes. False positives (FP) occur when the prediction is positive while the actual value is negative. Conversely, false negatives (FN) arise when the prediction is negative while the actual value is positive. By utilizing these counts, various performance evaluation measures are derived to assess the overall performance of the models.
@@
Table 5 reports the list of the metrics for the performance evaluation. A carefully selected set of statistical metrics is used to assess the performance from different perspectives. The metrics, Area Under the Receiver Operating Characteristic Curve (AUROC) and Area Under the Precision-Recall Curve (AUPRC), are employed to measure the models’ raw performance in the sense that these metrics do not require any threshold to convert probabilities into binary outcomes, making them suitable for evaluating the models’ pure performance. The Receiver Operating Characteristics (ROC) curve displays the entire set of possible combinations of true positive rates $TPR(c)=\frac{TP(c)}{\#\>Actual\>Positive}$ and false positive rates $FPR(c)=\frac{FP(c)}{\#\>Actual\>Negative}$ for some cutpoint $c\in(0,1)$ that maps the predicted probability to a binary category. Similarly, the Precision Recall (PR) curve plots the complete set of possible combinations of precision and recall for $c\in(0,1)$, where recall is a synonym for sensitivity. The Area under the ROC and PR curves increases with the underlying metrics for a given cutpoint. By aggregating over the entire set of cutpoints, these curves deliver a framework to assess the pure predictive ability of a forecasting model. Tharwat (2021) provides a comprehensive overview of other metrics and discusses their strengths and weaknesses, particularly in imbalanced binary classification problems.
@@ 4 Empirical results
The table reports the performance evaluation measures of forecasts obtained by logit models, ridge logit models, feed-forward (FFN), long short-term memory (LSTM), and gated recurrent unit (GRU) neural networks over different forecast horizons for the out-of-sample period, November 2006 to October 2021: Panel (A) presents the nowcasts. Panel (B), (C), (D), and (E) display the 1-month-ahead forecasts, the 3-month-ahead forecasts, the 6-month-ahead forecasts, and the 12-month-ahead forecasts, respectively.
@@
Table 6 presents the out-of-sample forecast performance of the models for various forecast horizons. In the nowcasting setup, there is a difference in predictive performance between the types of logit models and neural network models. The standard logit model does not outperform other models in any considered metric, although the difference is minor compared to other forecast horizons. The ridge logit model performs similarly to FFN. Although FFN performs slightly better, the differences are small for most metrics. However, LSTM and GRU perform significantly better in MCC and $F_{1}$-Score, mainly due to their high precision values. For instance, LSTM correctly identifies 90% of recession months, while 75% of its positive predictions are correct. In contrast, the ridge logit model has a precision of 51.5%, indicating that LSTM accurately predicts recessions without generating too many false alarms. The AUPRC values also highlight the superiority of LSTM and GRU models over the others, with GRU achieving an AUPRC of 0.837 compared to 0.642 for FFN and 0.529 for the ridge logit model. Davis and Goadrich (2006) argue that PR curves give a more informative picture of an algorithm’s performance when dealing with imbalanced datasets. The ratios of the number of months in booms to the number of months in recessions in 194 real-time vintages range from 4.5 to 6.7 with a mean value of 6, which means that the period of booms is, on average, six times longer than that of recessions. The percentage of recessions in the datasets fluctuates around 0.2 across the vintages. Although not highly skewed, the datasets are certainly imbalanced, and the difference in AUPRC values suggests that LSTM and GRU are better at extracting hidden patterns from scarce recession data.
@@
Moving to the immediate-term setup, logit models perform significantly worse, with MCC and $F_{1}$-Score approximately 40% lower than before. Conversely, the other models show similar or slightly improved performance. GRU exhibits the highest performance across various summarizing metrics such as balanced accuracy, MCC, and $F_{1}$-Score. Although the ridge logit model demonstrates better predictive performance in the short-term setup than LSTM, neural network models, particularly GRU, outperform the other models on average. GRU maintains a 90% accuracy in classifying recession months with higher specificity and precision, resulting in higher values of balanced accuracy, MCC, and $F_{1}$-Score. In the medium-term setup, LSTM performs slightly better than the other models, while GRU performs poorly, and thus, the difference between the two groups is minimal. However, LSTM achieves 90% sensitivity and over 90% specificity while maintaining precision above 0.5. The overall average performance of the models is the lowest among different forecast horizons. In the long-term setup, GRU surpasses the other models by a significant margin, with an $F_{1}$-Score of 0.737 compared to 0.468 for LSTM and 0.357 for the ridge logit model. GRU still accurately predicts 70% of recession months, with 77.8% of its positive predictions being correct.
@@
The figure depicts precision-recall (PR) curves for the out-of-sample period from November 2006 to October 2021: Panel (A) shows the PR curves for the nowcasting forecast horizon, comparing the logit model, ridge logit model, FFN, LSTM, and GRU. Panels (B), (C), (D), and (E) present the PR curves for the immediate-term (1-months-ahead), short-term (3-months-ahead), medium-term (6-months-ahead), and long-term (12-months-ahead) forecast horizon, respectively.
Neural network models, especially LSTM and GRU, significantly improve predictive performance compared to logit models. Across all metrics considered, neural network models consistently outperform logit models by a significant margin, except for the nowcasting setup. This superiority is evident in pure performance metrics, such as AUROC and AUPRC, and threshold-based point prediction metrics. The disparity is particularly prominent in AUPRC, a more reliable indicator when dealing with imbalanced data, highlighting the neural network models’ ability to handle class imbalance effectively. The performance difference decreases when comparing them to ridge logit models. Ridge logit models exhibit higher AUROC values than other model specifications, except for the long-term setup. However, regarding AUPRC, neural networks again prove to be a better choice. On average, neural network models perform better than ridge logit models. The gap widens further when considering only the two recurrent neural network models. Depending on the forecast horizons, LSTM or GRU may lead the competition, with occasionally substantial differences between them and the other models.
The ROC and PR curves of the models for five different forecast horizons are presented in Figures 6 and 7, respectively. Panel B and C of the ROC curves highlight the noticeable difference in forecast performance between the logit model (represented by the black line) and the other models. In Panel D and E, although the lines are closer together, there is still some discernible gap between each model. Overall, the ridge logit model, LSTM, and GRU exhibit superior performance of probability predictions across various forecast horizons. Figure 7 demonstrates that the green, blue, and purple lines lie, on average, above the red line and, even more so, above the black line. This finding supports the argument that, compared to linear models, neural network models are more capable of handling class imbalance in the data.
@@
The figure presents out-of-sample recession probabilities for the period from November 2006 to October 2021:: Panel (A) displays the predicted recession probabilities for the nowcasting forecast horizon, comparing the logit model, ridge logit model, FFN, LSTM, and GRU. Panels (B), (C), (D), and (E) show the predicted recession probabilities for the immediate-term (1-months-ahead), short-term (3-months-ahead), medium-term (6-months-ahead), and long-term (12-months-ahead) forecast horizon, respectively. The grey-shaded areas depict NBER recession months.
Figure 8 displays the out-of-sample forecasts in probabilities from November 2006 to September 2021, accompanied by the grey-shaded areas representing the Great Recession and the Covid-19 recession. Across all panels, similar patterns emerge: the graphs start at high levels before the Great Recession, either remain high or increase during that period, decrease afterward to a range between 0 and 0.25 for an extended period, and then exhibit significant fluctuations around the Covid-19 recession. The models successfully indicate economic and financial downturns during the Great Recession, but an external event like the COVID-19 recession remains largely unpredictable. Depending on the forecast horizon, once the model learns from the impact of the COVID-19 pandemic on the economy, it struggles to interpret this abrupt change and generates alternating forecasts.
@@ 5 Variable importance
To better understand the underlying factors driving the results, it is important to identify the main variables that have the most significant impact on the prediction. While linear models can refer to estimated coefficients or marginal effects to measure variable importance, more complex models like neural networks require different approaches. Molnar (2020) provides a comprehensive list of methods that enable the interpretation of complex machine-learning models. One such method is Local Interpretable Model-Agnostic Explanations (LIME), which suggests using other interpretable models to approximate the predictions locally. For an instance of interest, LIME constructs a new dataset by altering samples variable-wise, drawing from a normal distribution with sample mean and sample standard deviation, and obtains corresponding predictions from the black box model. This modified dataset is then used to train an interpretable model, such as Lasso or a decision tree, which is weighted by the proximity of the sampled instances to the instance of interest. The interpretable model approximates the black box model’s predictions at a local level, even though it may not accurately represent the global approximation.
@@
Another solution stems from cooperative game theory, particularly the Shapley value introduced by Shapley (1953). The Shapley value method involves assigning payouts to players based on their contribution to the overall payout. This cooperative framework resembles a game where players form coalitions and receive profits based on cooperation. In the context of model predictions and their interpretability, the game pertains to the prediction task, and the payout represents the difference between the actual prediction for that data point and the average prediction across all data points. The players correspond to the variable values of the data point, working together to achieve the payout, which means predicting a specific value. To calculate the Shapley value for a specific variable value, all possible coalitions of variable values, excluding the variable of interest, are formed for each data point. The variable values outside a coalition are substituted with random values of those variables from data to generate a prediction. The predictions are computed for each coalition, both with and without the variable value of interest. The difference between these predictions represents the marginal contribution of the variable value for that coalition. This computation is repeated for all possible coalitions, and the average of the marginal contributions across all coalitions yields the Shapley value for that variable value. Finally, the Shapley values for a variable can be averaged across data points to assess the relative importance of variables compared to each other.
@@
The Shapley value stands out as the sole explanation method supported by a robust theory that satisfies the axioms of efficiency, symmetry, dummy, and additivity to allow for fair distribution of the contributions among the variables (Molnar (2020)). Conversely, methods like LIME rely on the assumption of local linear behavior of the machine learning model but need a theoretical justification for why this approach is effective. Lundberg and Lee (2017) propose an alternative estimation method, the Shapley additive explanations (SHAP), which uses a kernel-based estimation approach for Shapley values (KernelSHAP). SHAP introduces an innovative aspect by presenting the Shapley value explanation as an additive variable attribution method, which can be viewed as a linear model. This perspective establishes a connection between LIME and Shapley values. In this context, SHAP defines the explanation model as follows:
@@
KernelSHAP is a method that estimates the contributions of individual variable values to the predictions by first generating a random coalition $c^{\prime}\in\{0,1\}^{K}$ consisting of K members by flipping a coin multiple times until we obtain a sequence of 0’s and 1’s. The sampled coalition of size $K$ is then used as a data point for the regression model. In this model, the target is the prediction for a coalition, whereby in the case of $c_{k}=0$, the absent variable value is substituted with random variable values from the data. This process is repeated for each data point. Then Shapley-compliant weights are computed according to the SHAP kernel proposed by Lundberg and Lee (2017). Finally, a weighted linear regression model is fitted on the modified data, and the model’s estimated coefficients are the Shapley values.
@@
The main distinction from LIME lies in how instances are weighted in the regression model. In LIME, the weighting is determined based on the proximity of instances to the original instance. The closer an instance is, the higher its weight in LIME. On the other hand, SHAP assigns weights to sampled instances based on the weight they would receive in the Shapley value estimation. Small coalitions (with fewer 1’s) and large coalitions (with many 1’s) receive the highest weights in SHAP. The underlying intuition is that the most knowledge about individual variables can be obtained when their effects can be studied in isolation. For a coalition consisting of a single variable, the isolated main effect of that variable on the prediction can be observed. When a coalition includes all variables except one, it reports the total effect of that particular variable, including the main effect and variable interactions. However, if a coalition comprises half the variables, it provides limited insight into the contribution of an individual variable due to the numerous possible coalitions with half of the variables.
@@
The SHAP values for logit models are computed using the KernelSHAP method using KernelExplainer from the SHAP package in Python. In contrast, the SHAP values for deep learning models are approximated by an enhanced version of the DeepLIFT algorithm introduced by Shrikumar et al. (2017) using DeepExplainer from the same package. DeepLIFT is a method used to analyze the output prediction of a neural network for a specific input. It achieves this by backpropagating the contributions of all units in the network to each input variable. DeepLIFT compares the activation of each unit to its reference activation, where for each variable, its sample mean is used as the reference input, and contribution scores are assigned based on the difference between them. Unlike other approaches, DeepLIFT can uncover dependencies that may be overlooked. It can also consider positive and negative contributions separately. Additionally, the scores can be efficiently computed in a single backward pass.
@@
SHAP method generates a matrix of Shapley values based on the number of data points and variables. These Shapley values can create global explanations by averaging the absolute Shapley values per variable across the twelve-time steps and data points. This process is repeated for the entire out-of-sample period. To compare the SHAP results of neural network models to linear models, GRU and the ridge logit model are selected for the subsequent analysis, as they perform better than the others in their respective groups of models. The results for GRUs are presented in Figure 9, which depicts the distribution of average absolute SHAP values across variables for the entire out-of-sample period. The variables are ranked based on their medians, providing insights into their importance for different forecast horizons.
@@
The ranking of variables may vary over time, but it offers an initial understanding of the key drivers influencing the prediction results for different forecast horizons. Variables associated with financial conditions such as the S&P 500 index and term spread, and macroeconomic variables related to GDP, inflation, and the housing market, commonly regarded as important recession indicators, consistently rank high (Estrella and Mishkin (1998)). These findings provide valuable insights into the black box and suggest that neural networks effectively capture the patterns inherent in the dynamics of the frequently cited recession indicators.
@@
Figure 10 presents the SHAP results for the ridge logit model. There are three major differences, among others, between the results in Figures 9 and 10. Firstly, the order of variables differs for each forecast horizon between the two types of models. While some variables appear in the leading groups for both models, their specific order varies significantly. For instance, the money supply M2 is influential in the ridge model but less prominent in the GRU model. Spearman’s rank correlations between them range from -0.24 to 0.10, indicating a weak correlation.
@@
This suggests that GRU and ridge models assign different weights to the variables. This difference could be attributed to the neural network’s ability to capture nonlinear relationships, allowing for varied weighting of the variables based on different patterns. Secondly, the SHAP values of the GRU models exhibit a more even distribution among the variables than the ridge model. This is reflected in the smaller variation in medians between the variables. The distinction becomes particularly pronounced in the short-term setting (up to 3 months), where the ridge model heavily depends on the S&P 500 index. Additionally, the ridge model exhibits numerous outliers in the SHAP values, indicating its high sensitivity to changes in macroeconomic and financial conditions in SHAP value estimation. In contrast, this highlights GRU as a more robust modeling framework. Finally, the within variation of the GRU-based SHAP values for some variables is significantly larger than others. Certain variables demonstrate little change in their SHAP values, while others exhibit significant variation. This phenomenon is less prominent in the outcomes of the ridge model. The variances are more uniformly spread, evident in the roughly equal sizes of boxes across variables. This is another indication of the GRU’s more resilient modeling framework. The GRU can adapt to shifts in economic conditions, adjusting the weights of specific variables accordingly. In contrast, the ridge model has relatively fixed variable orders with limited flexibility for variation.
The marginal effect analysis is another option to evaluate the variable importance of linear models. Since deep learning models can not be interpreted in terms of marginal effects, the average marginal effects of the ridge model are computed for the out-of-sample period and plotted in Figure 11. The Spearman’s rank correlations between the variable series for the ridge model, evaluated using either SHAP values or marginal effects, range from 0.22 to 0.51. The magnitudes of the correlations are relatively small. Still, the previous findings, especially the first one about the leading recession indicators, also apply to the ridge model evaluated by the marginal effects. The averaged LIME values in Figure 12, which may be regarded as the marginal effect version of GRU, reaffirm the main findings discussed above. The only minor distinction is that it emphasizes the term spread as the comprehensive recession indicator, both in short-term and long-term contexts.
@@
The SHAP method allows for a detailed analysis of the impact of a single variable on predictions. To anticipate recessions as early as possible, I focus on the long-term forecast horizon (12 months) and consider the term spread as the most relevant and influential variable for GRU to examine its effects on recession predictions for the GRU model. Figure 13 presents the dependence plot of the term spread using the latest available data from October 2021. The x-axis represents the standardized value of the variable (term spread), while the y-axis represents the corresponding SHAP value. Notably, there is a cutoff point around 0.2, above which a higher value of the term spread has a negative impact on the predicted recession probability. In contrast, a value below the cutoff positively influences the predicted probability. Additionally, the farther the value deviates from the cutoff, the stronger the effect of the term spread on the predicted probability becomes. This corroborates the traditional interpretation of the term spread about the recession probability according to Estrella and Mishkin (1996) and indicates that advanced neural network models like GRU uncover and utilize this relationship.